# Correlation-Embedded Transformer Tracking: A Single-Branch Framework

### v1

@@ Abstract
Developing robust and discriminative appearance models has been a long-standing research challenge in visual object tracking. In the prevalent Siamese-based paradigm, the features extracted by the Siamese-like networks are often insufficient to model the tracked targets and distractor objects, thereby hindering them from being robust and discriminative simultaneously. While most Siamese trackers focus on designing robust correlation operations, we propose a novel single-branch tracking framework inspired by the transformer. Unlike the Siamese-like feature extraction, our tracker deeply embeds cross-image feature correlation in multiple layers of the feature network. By extensively matching the features of the two images through multiple layers, it can suppress non-target features, resulting in target-aware feature extraction. The output features can be directly used for predicting target locations without additional correlation steps. Thus, we reformulate the two-branch Siamese tracking as a conceptually simple, fully transformer-based Single-Branch Tracking pipeline, dubbed SBT. After conducting an in-depth analysis of the SBT baseline, we summarize many effective design principles and propose an improved tracker dubbed SuperSBT. SuperSBT adopts a hierarchical architecture with a local modeling layer to enhance shallow-level features. A unified relation modeling is proposed to remove complex handcrafted layer pattern designs. SuperSBT is further improved by masked image modeling pre-training, integrating temporal modeling, and equipping with dedicated prediction heads. Thus, SuperSBT outperforms the SBT baseline by $4.7$%,$3.0$%, and $4.5$% AUC scores in LaSOT, TrackingNet, and GOT-10K. Notably, SuperSBT greatly raises the speed of SBT from $37$ FPS to $81$ FPS. Extensive experiments show that our method achieves superior results on eight VOT benchmarks. Code is available at https://github.com/phiphiphi31/SBT.
@@
1 Introduction
Visual object tracking (VOT) is a well-established subject in the field of computer vision. In recent years, there have been notable advancements in the visual tracking area. Nonetheless, it continues to pose a significant challenge, especially in real-world scenarios, due to the various factors that come into play, such as changes in lighting and object size, complex background clutter, and obstructions. Object tracking poses a significant challenge due to the fundamental yet competing goals it needs to achieve. On the one hand, it needs to recognize the target despite its changing appearance. On the other hand, it needs to filter out distractor objects in the background that may be very similar to the target.
@@ 1 Introduction
Figure 2: (a1) Standard Siamese-like feature extraction. (b1) Our single branch framework using joint feature extraction and correlation. Our pipeline removes separated correlation steps, e.g., Siamese cropping correlation [2], DCF [3] and transformer-based correlation [4]; (a2)/(b2) are the TSNE [5] visualizations of search features in (a1)/(b1) when feature networks go deeper.
@@
In the deep learning era, researchers have endeavored to address this challenge from two fundamental perspectives: enhancing the feature network and developing an elaborated correlation operation design. To achieve this, Siamese-like feature extraction, which utilizes deep convolutional neural networks (CNN) to learn more expressive feature embeddings, has significantly improved tracking performance. The recent emergence of vision transformer-based feature networks has further elevated the performance of the widely adopted Siamese-like feature extraction approach. On the other hand, developing a more robust correlation operation on top of the expressive feature embedding can help trackers differentiate targets from similar objects more easily. Typical Siamese trackers [6, 7] such as SiamRPN [2] adopt cross-correlation for efficiency while Discriminative Correlation Filter-based (DCF) trackers [8, 9] adopt online filter learning for discriminative power. Moreover, recent transformer-based fusion offers a solution for robust correlation design that balances both efficiency and accuracy.
@@
Existing popular tracking paradigms, i.e., Siamese and DCF-based trackers, can be formulated into three stages: Siamese-like feature extraction, correlation step, and final prediction step. Given that modern feature networks such as GoogLeNet [10] and ResNet [11] have become the predominant choice for feature extraction in the deep learning era, a significant portion of tracker research is dedicated to developing effective correlation operations that can differentiate targets from distractors based on their features. Despite their great success, few of these tracking paradigms notice the potential conflict between two competing goals, which can lead to a target-distractor dilemma and significant challenges during the correlation step. The underlying reasons can be divided into three categories: 1)1)1 ) The Siamese encoding process is unaware of the template and search images, leading to weaker instance-level discrimination of learned embeddings. 2)2)2 ) There is no explicit modeling for the backbone to learn the decision boundary that separates the two competing goals, leading to a sub-optimal embedding space. 3)3)3 ) During inference, arbitrary objects, including distractors can be tracked, whereas each training video only annotates one object. This gap is further widened by 2)2)2 ). Our key insight is that feature extraction should have dynamic instance-varying behaviors to generate “appropriate” embeddings for VOT to ease the target-distractor dilemma.
Based on these discussions, we propose a novel target-aware, fully transformer-based tracking framework built upon the attention scheme [12]. As shown in Fig. 2(a2), in contrast to the dual-branch Siamese feature extraction, our Single-Branch Transformer tracking (SBT) approach allows for deep interaction between the features of the two images during feature extraction. Therefore, extracting features depends on the target and asymmetry for an image pair, which enables the network to achieve a win-win scenario: it distinguishes the target from similar distractors while maintaining coherent characteristics among dissimilar targets. Our SBT framework can perform joint feature extraction and correlation within the vision backbone network. The effectiveness of the features obtained from SBT has been confirmed through the results presented in Fig. 2(d2). The features belonging to the target (green) are becoming increasingly distinct from the background (pink) and distractors (blue). In contrast, the search features from Siamese extraction are entirely unaware of the target.
@@
The overall framework of SBT is shown in Fig. 3. It has four critical architectural components, i.e., the Feature Relation Modelling (FRM) layer, the Patch Embedding (PaE) layer, the Positional Encoding (PE), and the prediction head network. The template and search images are split into image patches and then projected into feature tokens through the Patch Embedding layer. Feature tokens are fed into the stacked Feature Relation Modelling layer to fuse features within the same image or mix features across images layer by layer. FRM is a variant of the transformer layer in Visual Transformers [13, 14, 15] and there are two options of attention operator in the FRM layer, i.e., FRM-SA and FRM-CA, which are Self-Attention (SA) and Cross-Attention (CA), respectively. After joint feature extraction and correlation, the output feature tokens of the search image are directly fed to the prediction heads to obtain the location and size of the target. Our key technical innovation is the introduction of one single stream for template and search image pair processing that jointly extracts or correlates through homogeneous vision transformer layers.
Furthermore, we conduct extensive experiments on the vanilla SBT baseline to find the optimal network designs. Based on the insights, we summarize multiple general principles and propose an improved version, dubbed SuperSBT. Specifically, we first introduce a hierarchical structure to FRM layers to obtain the multi-scale feature representation. Then, we modify the FRM layers in shallow layers to enhance the local modeling ability of SBT. We further boost the tracking performance via a simple temporal modeling mechanism and Masked Image Modeling pre-training. SuperSBT obtains superior performance and outperforms SBT baseline, with even faster speed, as shown in Fig. 1. Both SBT and SuperSBT achieve promising results on eight VOT benchmarks while running at considerable speed.
@@
This study builds upon our two conference papers (DualTFR [16] and SBT [17]) and significantly extends them in various aspects. For example, we develop an improved tracker dubbed SuperSBT, greatly enhancing the SBT baseline. The new contributions are summarized as follows.
@@
  • We propose more specialized architectural modules for tracking, which greatly raises the overall tracking performance of our SBT baseline and speeds up online tracking.
@@
  • We have conducted experiments on more benchmarks, e.g., TrackingNet [18], TNL2k [19], and VOT2021 [20]. We have also included more comparisons with recent strong trackers.
@@
  • We have provided more thorough discussions and details.
2 Related Work
@@ 2.2 Vision Transformer
2.3 Transformer in Tracking
@@
Figure 3: (a) Architecture of our proposed Single-Branch Transformer framework for tracking (SBT). Unlike Siamese, DCF, and Transformer-based methods, it has no standalone module for computing correlation. Instead, it embeds correlation in all Feature Relation Modelling (FRM) layers at different network levels. The fully fused features of the search image are directly fed to the prediction network to obtain the localization and size of the target. (b) shows the structure of the FRM layer, which is a variant of the transformer [13] layer. There are two options for attention operators in the FRM layer, i.e., Self-Attention (SA) and Cross-Attention (CA). SA operator fuses features within the same image while the CA operator mixes features across images.
3 Single-Branch Transformer Tracking
@@ 3.1 Overall Architecture of SBT
3.2 Transformer-Based Backbone
@@
Patch embedding. The patch embedding layer is to embed the two images into feature tokens through a convolutional layer. The kernel size and stride of the convolutional layer in the patch embedding layer can be modified for different model structures, and we conduct exploration studies in Sec. 3.4.
Table I: The left part compares different factors of SBT, including attention computation methods (ATTN), position encoding methods (PE), patch embedding methods (PaE), number of model parameters, and flops. The right part compares the rest of the factors based on $\rm A_{5}$ (described in the left part), such as the feature dimensions (DIM) and the number of blocks (BLK), as well as the stride of the feature maps in each stage. All models, unless explained, follow the same setting: training from scratch, interleaved FRM-SA/FRM-CA block in the third stage, $128$ for template image and $256$ search image. All the experiments follow the official GOT-10k [1] test protocol.
@@
Feature relation modeling. The Feature Relation Modeling (FRM) layer can simultaneously model cross-image and intra-image relationships. As shown in Fig 3(b), FRM layer follows the structure of a standard transformer layer [12, 13]. Intuitively, the FRM layer gradually fuses features from the same and different images through Self-Attention (SA) and Cross-Attention (CA), respectively. In each FRM layer, the template $f_{z}$ and search features $f_{x}$ are reshaped to feature tokens and then embedded into the Query/Key/Value space. Let $\chi_{(.)}$ denote a function that reshapes/arranges feature maps into the desired form. The function varies for different methods. We compute the $q,k,v$ features as:
@@
where $\{\omega_{q},\omega_{k},\omega_{v}\}$ represent linear projections. Many works [15, 14] attempt to modify the attention operators with task priors. For example, the Vanilla Global attention (VG) [13] computes attention among all feature tokens. So $\{\chi_{q},\chi_{k},\chi_{v}\}$ represent identity mapping. The Spatial-Reduction Global attention (SRG) [15, 53] uses a convolution with a stride larger than one (i.e., $\{\chi_{k},\chi_{v}\}$) to reduce the spatial resolution of the key and value features. The resolution of the query features is not changed. Then, it computes global attention as VG. The Vanilla Local window attention (VL) [54] split feature tokens in groups based on their spatial locations and only compute attention within each group. Swin Transformer [14] further adds a Shift window mechanism to vanilla Local attention (SL) for global modeling. More empirical studies and discussions are in Sec. 3.4. Here, we omit the various attention operators in the formula for clarity. The following equation shows how we compute self-attention and cross-attention to model intra-image and cross-image relationships:
@@
Position encoding. Since template and search features are reshaped into feature tokens, Positional Encoding (PE) embeds spatial information into sequential tokens. Non-parametric or conditional methods can implement positional encoding. For majority methods [13, 14, 40], the encoding is generated by the sinusoidal functions with Absolute coordinates (Abs) or Relative distances (Rel) between tokens. Being much simpler, Conditional positional encoding [56, 15, 53] (Cond) generates dynamic encoding by convolutional layers. In our SBT model, we add a $3\times 3$ depth-wise convolutional layer $\varphi_{pe}$ to MLP before GELU as conditional positional encoding. We further improve it by using relative positional encoding. More details are presented in Sec. 3.4.
@@
3.3 Direct Prediction Head
Differing from existing tracking methods, we directly attach a classification head and regression head onto the search feature from the SBT feature network without any additional correlation operations. In this work, we adopt a lightweight convolutional-based anchor-free head [57, 27, 46], as illustrated in Sec. 5.
@@ 3.3 Direct Prediction Head
3.4 Exploration Studies on SBT
In this section, we analyze the impact of different structure designs in SBT tracking and assess their performance. The effective design principles are summarized below, which can be used to enhance the SBT tracking pipeline in Sec. 4.
@@ 3.4 Exploration Studies on SBT
Hierarchical transformer structure is more effective than plain structure. Based on the findings presented in Tab. I, it is evident that the model variants from $\rm A_{2}$ to $\rm A_{7}$ outperform the single-stage model variant $A_{1}$, despite having a similar model size. This suggests that a hierarchical structure is more effective due to its ability to provide a multi-scale representation.
@@
Layer pattern in the shallow and deep stage can be designed differently. The optimal placement pattern significantly impacts the performance of FRM-CA layers. So, we attempt to place the FRM-SA/FRM-CA layer differently. In Fig. 4(f), we surprisingly find that the interleaved FRM-SA/FRM-CA pipeline performs better than the separation pattern even with less Cross-Attention and latter earliest cross position ($\rm C_{3}$ vs. $\rm C_{1}$). The potential cause is that the FRM-SA layer can refine the template/search feature after the correlation, resulting in a more expressive feature space for matching. In Fig. 4(f), model ($\rm C_{9}$) achieves the best performance $67.2\%$ when the interval is $1$. When the interval increases to 2, the performance drops from $61.1\%$ to $59.2\%$ ($\rm C_{3}$ vs. $\rm C_{10}$). From Fig. 4(b) and Fig. 4(c), we observe that early-cross in shallow-level (stage 1 and 2) does not bring many improvements ($\rm C_{2}$ vs. $\rm C_{8}$, $\rm C_{6}$ vs. $\rm C_{7}$) but lowers the inference speed. This is because initiating the cross too early disrupts the one-shot inference, and the shallow feature is less expressive.
@@
4 Improved Single-Branch Tracking

### v2

@@ Abstract
Developing robust and discriminative appearance models has been a long-standing research challenge in visual object tracking. In the prevalent Siamese-based paradigm, the features extracted by the Siamese-like networks are often insufficient to model the tracked targets and distractor objects, thereby hindering them from being robust and discriminative simultaneously. While most Siamese trackers focus on designing robust correlation operations, we propose a novel single-branch tracking framework inspired by the transformer. Unlike the Siamese-like feature extraction, our tracker deeply embeds cross-image feature correlation in multiple layers of the feature network. By extensively matching the features of the two images through multiple layers, it can suppress non-target features, resulting in target-aware feature extraction. The output features can be directly used to predict target locations without additional correlation steps. Thus, we reformulate the two-branch Siamese tracking as a conceptually simple, fully transformer-based Single-Branch Tracking pipeline, dubbed SBT. After conducting an in-depth analysis of the SBT baseline, we summarize many effective design principles and propose an improved tracker dubbed SuperSBT. SuperSBT adopts a hierarchical architecture with a local modeling layer to enhance shallow-level features. A unified relation modeling is proposed to remove complex handcrafted layer pattern designs. SuperSBT is further improved by masked image modeling pre-training, integrating temporal modeling, and equipping with dedicated prediction heads. Thus, SuperSBT outperforms the SBT baseline by $4.7$%,$3.0$%, and $4.5$% AUC scores in LaSOT, TrackingNet, and GOT-10K. Notably, SuperSBT greatly raises the speed of SBT from $37$ FPS to $81$ FPS. Extensive experiments show that our method achieves superior results on eight VOT benchmarks.
@@
I Introduction
Visual object tracking (VOT) is a well-established subject in the field of computer vision. In recent years, there have been notable advancements in the visual tracking area. Nonetheless, it continues to pose a significant challenge, especially in real-world scenarios, due to various factors, such as changes in lighting and object size, complex background clutter, and obstructions. Object tracking poses a significant challenge due to the fundamental yet competing goals it needs to achieve. On the one hand, it needs to recognize the target despite its changing appearance. On the other hand, it needs to filter out distractor objects in the background that may be very similar to the target.
@@ I Introduction
Figure 2: (a1) Standard Siamese-like feature extraction. (b1) Our single-branch framework uses joint feature extraction and correlation. Our pipeline removes separated correlation steps, e.g., Siamese cropping correlation [2], DCF [3] and transformer-based correlation [4]; (a2)/(b2) are the TSNE [5] visualizations of search features in (a1)/(b1) when feature networks go deeper.
@@
In the deep learning era, researchers have endeavored to address this challenge from two fundamental perspectives: enhancing the feature network and developing an elaborated correlation operation design. To achieve this, Siamese-like feature extraction, which utilizes deep convolutional neural networks (CNN) to learn more expressive feature embeddings, has significantly improved tracking performance. The recent emergence of transformer-based feature networks has further elevated the performance. On the other hand, developing a more robust correlation operation can help trackers differentiate targets from similar objects more easily. Typical Siamese trackers [6, 7] such as SiamRPN [2] adopt cross-correlation for efficiency while Discriminative Correlation Filter-based (DCF) trackers [8, 9] adopt online filter learning.
@@
Given that modern CNN-based visual backbones [10, 11, 12] have become the predominant choice for feature extraction, most trackers are dedicated to developing effective correlation operations that can differentiate targets from distractors based on their features. Despite their great success, we infer that Siamese-like tracking paradigms are trapped by the target-distractor dilemma caused by the two competing goals mentioned. The underlying reasons are as follows: 1)1)1 ) The Siamese encoding process is unaware of the template and search images, leading to weaker instance-level discrimination of learned embeddings. 2)2)2 ) There is no explicit modeling for the backbone to learn the decision boundary that separates the two competing goals, leading to a sub-optimal embedding space. 3)3)3 ) During inference, arbitrary objects, including distractors, can be tracked, whereas each training video only annotates one object. This gap is further widened by 2)2)2 ). Our key insight is that feature extraction should have dynamic instance-varying behaviors to generate appropriate embeddings for VOT to ease the target-distractor dilemma.
@@
To address these issues, we propose a novel single-branch, fully transformer-based tracking framework that allows for deep interaction between the features of the two images during feature extraction. In contrast to the dual-branch Siamese pipeline in Fig. 2(a1), our Single-Branch Transformer tracking (SBT) approach achieves a win-win scenario at the stage of feature extraction, i.e., it distinguishes the target from similar distractors while maintaining coherent characteristics among dissimilar targets. The effectiveness of the features obtained from SBT has been confirmed through the results presented in Fig. 2(d2). The overall framework of SBT is shown in Fig. 3. It has four critical architectural components: the Feature Relationship Modeling (FRM) layer, the Patch Embedding (PaE) layer, the Positional Encoding (PE) layer, and the prediction head. The template and search images are split into image patches and then projected into feature tokens through the patch embedding layer. Feature tokens are fed into the stacked feature relation modeling layer to fuse features within the same image or mix features across images layer by layer. FRM is a variant of the transformer layer in vision transformers [13, 14, 15]. The search image features obtained from SBT are directly fed to the prediction head. Our key technical innovation is the single stream for image pair feature processing that jointly extracts and correlates through homogeneous transformer layers.
To better adapt the single-branch pipeline into tracking, we conduct extensive experiments to summarize design principles and propose an improved SBT version, dubbed SuperSBT. Specifically, we first introduce a three-stage hierarchical structure to obtain a multi-scale feature representation with a favorable feature downsampling ratio of $16$ in tracking. Then, we decouple the layer pattern design in the shallow and deep stages to obtain a more expressive feature embedding for tracking. In the deep stage, we design a unified relation modeling layer as the main building unit, which is friendly to cross-image modeling and parallel computing. The unified relation modeling layer can implement intra-image and cross-image relation modeling by a global attention scheme. In the shallow stage, we propose a local modeling layer to enhance the local modeling ability, which replaces the attention scheme with convolutional filters. We also choose the optimal network variants based on the design principles, balancing speed, performance, and model size. Three variants of SuperSBT (light, small, and base versions) are proposed and obtain promising results in terms of speed and performance. We further explore the temporal modeling ability of SuperSBT via a simple dynamic template scheme, which concatenates an extra template at the beginning of input. Moreover, our SuperSBT network is pre-trained via Masked Image Modeling (MIM) on abundant unpaired images such as ImageNet [16], leading to a fast convergence and a better tracking performance in fine-tuning. In contrast to the SBT baseline, the SuperSBT, with optimal network designs, obtains superior performance on eight VOT benchmarks, with even faster speed, as shown in Fig. 1.
@@
This study builds upon our conference paper SBT [17] and is one step forward for our conference paper DualTFR [18]. It significantly extends them in various aspects. For example, we develop an improved tracker dubbed SuperSBT, greatly enhancing the SBT baseline. The new contributions are summarized as follows.
@@
  • We propose an improved version dubbed SuperSBT, which has more specialized architectural adaptations for visual tracking, which greatly raises the overall tracking performance of our SBT baseline and speeds up online inference.
@@
  • We have proposed more model variants and conducted experiments on more visual tracking benchmarks, e.g., TrackingNet [19], TNL2k [20], and VOT2021 [21]. We have also included more comparisons with recent strong trackers.
@@
  • We have provided more thorough design principles, discussions, and details.
II Related Work
@@ II-B Vision Transformer
Figure 3: (a) Architecture of our proposed Single-Branch Transformer framework for tracking (SBT). Unlike Siamese, DCF, and Transformer-based methods, it has no standalone module for computing correlation. Instead, it embeds correlation in all Feature Relation Modeling (FRM) layers at different network levels. The fully fused features of the search image are directly fed to the prediction network to obtain the localization and size of the target. (b) shows the structure of the FRM layer, which is a variant of the transformer [13] layer. There are two options for attention operators in the FRM layer, i.e., Self-Attention (SA) and Cross-Attention (CA). SA operator fuses features within the same image while the CA operator mixes features across images.
II-C Transformer in Tracking
@@
III Single-Branch Transformer Tracking
@@ III-A Overall Architecture of SBT
III-B Transformer-Based Backbone
@@
Patch embedding. The patch embedding layer is to embed the two images into feature tokens through a convolutional layer. The kernel size and stride of the convolutional layer in the patch embedding layer can be modified for different model structures, and we conduct exploration studies in Sec. III-D.
@@
Table I: The left part compares different factors of SBT, including attention computation methods (ATTN), position encoding methods (PE), patch embedding methods (PaE), number of model parameters, and flops. The right part compares the rest of the factors based on $\rm A_{5}$ (described in the left part), such as the feature dimensions (DIM) and the number of blocks (BLK), as well as the stride of the feature maps in each stage. All models, unless explained, follow the same setting: training from scratch, interleaved FRM-SA/FRM-CA block in the third stage, $128$ for template image, and $256$ search image. All the experiments follow the official GOT-10k [1] test protocol.
Feature relation modeling. The Feature Relation Modeling (FRM) layer can simultaneously model cross-image and intra-image relationships. As shown in Fig 3(b), the FRM layer follows the structure of a standard transformer layer [40, 13]. Intuitively, the FRM layer gradually fuses features from the same and different images through Self-Attention (SA) and Cross-Attention (CA), respectively. In each FRM layer, the template $f_{z}$ and search features $f_{x}$ are reshaped to feature tokens and then embedded into the Query/Key/Value space. Let $\chi_{(.)}$ denote a function that reshapes/arranges feature maps into the desired form. The function varies for different methods. We compute the $q,k,v$ features as:
@@
where $\{\omega_{q},\omega_{k},\omega_{v}\}$ represent linear projections. Many works [15, 14] attempt to modify the attention operators with task priors. For example, the Vanilla Global attention (VG) [13] computes attention among all feature tokens. So $\{\chi_{q},\chi_{k},\chi_{v}\}$ represent identity mapping. The Spatial-Reduction Global attention (SRG) [15, 54] uses a convolution with a stride larger than one (i.e., $\{\chi_{k},\chi_{v}\}$) to reduce the spatial resolution of the key and value features. The resolution of the query features has not been changed. Then, it computes global attention as VG. The Vanilla Local window attention (VL) [55] split feature tokens in groups based on their spatial locations and only compute attention within each group. Swin Transformer [14] further adds a Shift window mechanism to vanilla Local attention (SL) for global modeling. More empirical studies and discussions are in Sec. III-D. Here, we omit the various attention operators in the formula for clarity. The following equation shows how we compute self-attention and cross-attention to model intra-image and cross-image relationships:
@@
Position encoding. Since template and search features are reshaped into feature tokens, Positional Encoding (PE) embeds spatial information into sequential tokens. Non-parametric or conditional methods can implement positional encoding. For majority methods [13, 14, 41], the encoding is generated by the sinusoidal functions with Absolute coordinates (Abs) or Relative distances (Rel) between tokens. Being much simpler, Conditional positional encoding [57, 15, 54] (Cond) generates dynamic encoding by convolutional layers. In our SBT model, we add a $3\times 3$ depth-wise convolutional layer $\varphi_{pe}$ to MLP before GELU as conditional positional encoding. We further improve it by using relative positional encoding. More details are presented in Sec. III-D.
@@
III-C Direct Prediction Head
Differing from existing tracking methods, we directly attach a classification head and regression head to the search feature from the SBT feature network without any additional correlation operations. In this work, we adopt a lightweight convolutional-based anchor-free head [58, 28, 47], as illustrated in Sec. V.
@@ III-C Direct Prediction Head
III-D Exploration Studies on SBT
In this section, we analyze the impact of different structure designs in SBT tracking and assess their performance. The effective design principles are summarized below, which can be used to enhance the SBT tracking pipeline in Sec. IV.
@@ III-D Exploration Studies on SBT
Hierarchical transformer structure is more effective than plain structure. Based on the findings presented in Tab. I, it is evident that the model variants from $\rm A_{2}$ to $\rm A_{7}$ outperform the single-stage model variant $A_{1}$, despite owning a similar model size. This suggests that a hierarchical structure is more effective due to its ability to provide a multi-scale representation.
@@
Layer patterns in the shallow and deep stages can be designed differently. The optimal placement pattern significantly impacts the performance of FRM-CA layers. So, we attempt to place the FRM-SA/FRM-CA layer differently. In Fig. 4(f), we surprisingly find that the interleaved FRM-SA/FRM-CA pipeline performs better than the separation pattern even with less Cross-Attention and latter earliest cross position ($\rm C_{3}$ vs. $\rm C_{1}$). The potential cause is that the FRM-SA layer can refine the template/search feature after the correlation, resulting in a more expressive feature space for matching. In Fig. 4(f), model ($\rm C_{9}$) achieves the best performance $67.2\%$ when the interval is $1$. When the interval increases to 2, the performance drops from $61.1\%$ to $59.2\%$ ($\rm C_{3}$ vs. $\rm C_{10}$). From Fig. 4(b) and Fig. 4(c), we observe that early-cross in shallow-level (stage 1 and 2) does not bring many improvements ($\rm C_{2}$ vs. $\rm C_{8}$, $\rm C_{6}$ vs. $\rm C_{7}$) but lowers the inference speed. This is because initiating the cross too early disrupts the one-shot inference, and the shallow feature is less expressive.
@@
IV Improved Single-Branch Tracking